In a process with a regulated mass system, there is a heat loss (heat exiting the system) of 1055 J with the energy in the system of 210 kJ. Define the work being done and include compression or expansion?
1
Expert's answer
2020-10-13T18:02:26-0400

given heat loss of 1055J , Q = - 1055J

energy in the system, increase in energy "\Delta U = 210 KJ"   

net work done

"W = Q - \Delta U = -1055 - 210000 = - 211055 J"

here work done is negative so work is done on the system and it is compression process.

Now if the given heat loss is in KJ, i mean Q = -1055 KJ

then net work done

"W = Q - \Delta U = -1055 - 210= - 1265KJ"

here also work done is negative so work is done on the system and it is compression process